# Measuring Star Formation Histories from Asymptotic Giant Branch Stars: A Demonstration in M31

### v2

@@ 4 Data
From the raw source catalogs, we keep only sources classified as stellar in both J and K by the CASU pipeline. We then clean the final catalogs using photometric quality cuts that are demonstrated in the panels of Figure B1. We use a constantexponential function to cull the data via the the photometric uncertainty as a function of the J-band and K-band magnitudes. The specific functional form of these cuts can be found in Appendix B.
We generate ASTs to characterize the biases, uncertainties, and completeness of our data using the procedure described in Ren et al. (2024). We run ASTs across the PHAT footprint. However, because of the wide areal coverage of the halo data (), running ASTs for each position of the UKIRT footprint is prohibitively computationally expensive in the framework of this photometric routine. Instead, we run ASTs in five representative stellar density ranges, using stellar density as an indicator of the expected crowding (and resulting photometric noise) in each range. We then measure the average stellar density of the 18 halo regions discussed in §5.3. We use the set of ASTs that were generated for regions of similar stellar density. Table 1 lists the properties of the ASTs used. The process of generating ASTs in select regions of specific stellar density and applying them more broadly is also used by the PHAT/PHATTER teams, who faced a similar computational challenge of running ASTs across the entire extent of their M31 and M33 footprints, respectively (Williams et al., 2017; Lazzarini et al., 2022). We summarize our ASTs in Table 1.
@@ 5.1 Application to Mock Data
To measure the AGB star SFH, we run MATCH on the ‘observed’ CMDs shown in Figure 4. We adopt an age range of to with a spacing of 0.1 dex and a metallicity range of [M/H] = -0.4 to 0.0 dex with a spacing of 0.1 dex. We use well-populated basis functions () to minimize any effects of Poisson variations, which are the not purpose of this test.
Next, we run the simulated CMDs through MATCH as if they were real photometric datasets. We then compute random uncertainties using HMC. We ran an exhaustive set of simulations with various SFHs. For brevity, we only explore a couple of illustrative examples below for a constant and bursty SFH. However, the conclusions are the same no matter in the input SFH: MATCH self-consistency recovers the input SFH using mock data.
Synthetic Galaxy I: Multi-burst SFH. Synthetic galaxy I was created from three star formation bursts with ages of 2.0, 5.0, 7.9 Gyr and metallicities [M/H] = -0.4, -0.2, 0.0 dex, respectively. Each burst forms a stellar mass of . In the top panel of Figure 4, we show the “observed” CMD, recovered best fit model CMD, and residual CMD, along with the recovered and input star formation histories.
@@
The excellent agreement between all input and recovered tests indicates that MATCH is self-consistently able to handle our custom basis functions. We now turn to application of this method to real data.
@@ 5.2 Comparison to PHAT
A further check on consistency is to compare our AGB star-based SFHs to those derived from deep optical CMDs. Ideally, we would use an optical CMD that reaches the oldest MSTO to provide an anchor for this testing.
However, only datasets that are suitably deep, outside the Magellanic Clouds that have been used to calibrate the AGB models, are from small-area HST or JWST fields (e.g., Skillman et al. 2003; Dolphin et al. 2003; Brown et al. 2006; Geha et al. 2015; Albers et al. 2019; McQuinn et al. 2024a) or from ground-based imaging of MW satellites (e.g., de Boer et al. 2011, 2012a, 2012a, 2014; Muñoz et al. 2018). At issue is that these fields do not contain a suitable number of AGB stars (at least , as we discuss below) to enable a reliable SFH.
We instead opt to compare our AGB star-based SFH to those measured as part of the PHAT survey. Specifically, we use the SFH fits from W17, who measured the spatially resolved ancient SFH of the PHAT region from optical CMDs that extend to mag below the RC, as shown in Figure 5. Although these CMDs do not reach the 10 Gyr MSTO, and instead rely on advanced phases of evolution (e.g., the position and shape of the RC), W17 took great care to validate their results against CMDs taken throughout M31’s disk that did reach the oldest MSTO (Bernard et al., 2015a, b). We did consider comparing results directly to other M31 SFHs that do reach the oldest MSTO (Brown et al., 2006; Bernard et al., 2015a), but there are not enough AGB stars in the spatially matched HST footprints. Similarly, we considered a comparison to the larger number of fields from Richardson et al. (2008), but (a) they do not provide quantitative SFHs and (b) we are again limited by the number of AGB stars in a given field.
Recognizing the limitations of their moderately deep CMDs, W17 adopt a single ancient age bin of 8-14 Gyr ago. They adopt a finer time resolution (0.1 dex) for ages younger than Gyr. Therefore, for comparison purposes, in this paper we only measure the AGB-based SFHs back to 8 Gyr ago. We stress this age limit is not inherent to the AGB star SFH method itself, but rather due to our desire for empirical validation.
@@ 5.2.1 Photometric Selection
Beyond an age restriction, we limit this comparison to the outer region ( kpc) of the PHAT survey, as illustrated in Figure 6. This enables us to compare to the deepest optical data available, avoids the need to model the complex spatially varying differential extinction that dominates most of M31’s inner disk (e.g., Dalcanton et al., 2015; Lewis et al., 2015). It also mitigates the impact of crowding on the optical CMDs and hence the SFHs. Crowding significantly impacts the depth of the optical CMDs in M31’s inner disk, reducing the fidelity of the W17 SFHs. We note that while crowding affects the faintest stars in the the PHAT survey, it has minimal impact on brighter stars. For example, Figure 6 in Gregersen et al. (2015) demonstrates the RGB stars with kpc have completeness.
@@
In comparison, our AGB star-based SFH contains AGB stars that are located within the selection box in Figure 5, which is defined on the faint end by the 50% completeness limit of our data ( mag). Unlike the the PHAT optical CMDs, our AGB star data are not generally crowded in any part of M31. The 50% completeness limit in the inner regions ( kpc) is only mag brighter than in the outskirts. The significantly reduced sensitivity to crowding is another strength of AGB stars in the NIR for SFH determinations.
@@
We opt to use the Draine et al. (2014) dust map, which is based on far-infrared and submillimeter emission observations from the Spitzer Space Telescope and Herschel Space Observatory. The Dalcanton et al. (2015) reddening map predicts less extinction than the Draine et al. (2014) dust map, although their overall morphological agreement is excellent. Thus, we use the Draine et al. (2014) map with the Dalcanton et al. (2015) corrections. Specifically, the Dalcanton et al. (2015) corrects the Draine et al. (2014) maps by dividing by a factor of and then an additional additive correction of to account for possible systematic biases. This transformation is then as follows: .
@@
Figure 5 shows a comparison between the HST-based optical CMD of M31 at kpc (left) and our ground-based NIR CMD of the same region (right). As discussed in §2, it is virtually impossible to separate out AGB and RGB stars in the optical. This point is reinforced by the optical PHAT CMD.
In the NIR, we select AGB stars as having colors between mag and magnitudes between mag based on the locations AGB stars isochrones. This selection also eliminates contamination from red supergiants and foreground stars, both of which are bluer than our selection region. We also limit the AGB stars as having magnitudes brighter than the TRGB ( mag), which is also the 50% completeness limit of our data.
In total, there are million stars in the PHAT region above a magnitude of mag (the 50% completeness limit) used to measure the SFH, compared to just AGB stars used to measure our SFH. We emphasize that most of the age information in the PHAT SFHs is from the position of the red clump at mag, whereas we only used AGB stars brighter than mag in NIR wavelengths.
5.2.2 Measuring the AGB Star-based SFH in the PHAT Region
To measure the AGB star SFH, we run MATCH on stars within the grey box in the UKIRT CMD shown in Figure 5. We adopted an age range of to with a spacing of 0.1 dex and a metallicity range of [M/H] to dex also with a spacing of 0.1 dex. The metallicity range was motivated by measurements of the RGB star metallicities across M31’s disk by the PHAT team (Gregersen et al., 2015), while the age range largely matches that of W17 and the age range permitted by AGB stars.
@@ 5.2.2 Measuring the AGB Star-based SFH in the PHAT Region
For our fitting, we adopt an M31 distance modulus of mag ( kpc; Lee 2023). This distance was derived from the J-region asymptotic giant branch (JAGB) method, a standard candle based on AGB stars, using the same UKIRT photometry studied in this paper. This is virtually identical to the distance modulus assumed by W17 and to other commonly adopted distances to M31 (e.g., Jeffery et al., 2011; de Grijs & Bono, 2014; Bhardwaj et al., 2016; Savino et al., 2022). We use a CMD bin size of mag in color and magnitude, respectively. We then follow the CMD modeling procedure described in §3.
@@
Figure 7 shows the SFH of the PHAT region based on AGB stars. From the residual CMDs (left panel), we see that our best fit model provides a good match to the data. This is visually apparent by the general ‘checkerboard’ pattern in the residual significance plot, i.e., the residual weighed in each Hess diagram in units of standard deviations. This plot shows modest discrepancies between the data and model at the bluest colors, but they are of moderate significance and are on-par with the quality of residuals in most optical CMD modeling (e.g., Monelli et al. 2010; Hidalgo et al. 2011; Weisz et al. 2011, 2014; Skillman et al. 2017; Albers et al. 2019; Savino et al. 2023) and with JWST in the NIR (e.g., McQuinn et al., 2024a, c).
@@
The middle panel shows the cumulative SFH the AGB star-based SFH in black, and the W17 best fit PARSEC-based SFH in red. The uncertainty envelope for both SFHs reflects the 68% random uncertainties. We note that although W17 include AGB stars in their CMD modeling they (a) are of limited utility due to overlap with the RGB and (b) are not fit with the more recent versions of the COLIBRI AGB models, as they were not yet available. As discussed extensively in W17, the age information at most epochs comes from the RC and ratios of RC to RGB stars, with some contribution from AGB stars. The two SFHs, (AGB star and W17) use the same IMF, but adopt slightly different minumum mass limits for the low-mass IMF. This introduces an order of unity difference in the SFHs and stellar masses formed. We apply this small IMF correction to the W17 for the purposes of equal comparison.
@@
The two cumulative SFHs are in excellent agreement. Both show a modest amount of stellar mass formed from Gyr ago, followed by two episodes of enhanced star formation. THe AGB SFH has a burst Gyr ago, while the optical SFH shows a less intense burst from Gyr ago. Both formed the same stellar mass over the 3-5 Gyr ago interval. Though the difference in timing is not statistically significant, in future studies it will be interesting to understand if such differences could be due to information content of the data (e.g., ‘bleeding’ of star formation into adjacent time bins due to the increased age-metallicity degeneracy in the RC). The second star formation event is from Gyr ago (AGB stars) or Gyr ago in the optical. The total stellar mass formed over 8 Gyr by each method is the same.
Random uncertainties on the W17 SFH are larger than for the AGB stars, despite a factor of more stars used to measure the optical CMD-based SFH. This is most likely due to the reduced age-metallicity-dust degeneracy of the AGB stars in the NIR compared to the combination of the MS, RC, RGB, and AGB stars in the optical.
@@
We further investigated this by comparing the age-metallicity relationships (AMRs) from our fits and W17. We find the AMRs to be nearly identical between the two methods. Both are generally consistent with Solar metallicity, except for a noticeable drop to [M/H] at Gyr. Similar decreases in metallicity are measured in other regions of M31’s disk (Bernard et al., 2015a) from CMDs that reach the oldest MSTO . This provides additional confidence in the fidelity of our AGB star-only SFH.
@@
The right panel of Figure 7 shows the differential SFH, i.e., SFR vs time, of the AGB star and PHAT SFHs. As with the cumulative SFHs, we find good agreement between the amplitude and timing of star formation in both SFHs, though there are a few differences. For example, the timing and amplitude of star formation around Gyr ago and Gyr ago are slightly different. As with the cumulative SFHs, such differences in the exact timing of star formation from CMD modeling are often observed when using different stellar models or CMD fitting codes (e.g., Weisz et al., 2011; Dolphin, 2012; Cole et al., 2014; Skillman et al., 2017; Savino et al., 2023).
@@
To examine how significant these variations may be, in Figure 8 we compare best fit SFHs measured using different stellar models from W17 with our AGB star-based SFH. It is clear from the cumulative and differential SFHs that our AGB star-only SFH is well within the model-to-model variance from optical CMD analysis. Each of the optical-based SFHs shows the same qualitative trend of stellar mass growth, but with slightly different timings and amplitudes of star formation episodes, as well as modest differences in the total stellar mass formed. Within this context, our AGB-based SFHs are as robust as any of the SFHs based on deeper optical CMDs.
@@
Our measurement is based on AGB stars. In §A, we quantify how the quality of the fit and statistical errors scale with the number of AGB stars, by measuring the AGB SFH from decreasingly smaller random samples of AGB stars. In Figure 16, we show that as few as 280 AGB stars can deliver a reasonable SFH compared to the W17.
Finally, we note that our NIR CMD of the PHAT region (Figure 5) clearly contains AGB stars that are older than 8 Gyr. While in principle we could extend our analysis to older ages (as discussed in Appendix C), for the sake of empirical verification, we limit this comparison to 8 Gyr and will expand our tests to older look back times in future papers, when suuitable datasets are available.

### v3

@@ 4 Data
From the raw source catalogs, we keep only sources classified as stellar in both J and K by the CASU pipeline. We then clean the final catalogs using photometric quality cuts that are demonstrated in the panels of Figure A1. We use a constantexponential function to cull the data via the the photometric uncertainty as a function of the J-band and K-band magnitudes. The specific functional form of these cuts can be found in Appendix A.
We generate ASTs to characterize the biases, uncertainties, and completeness of our data using the procedure described in Ren et al. (2024). We run ASTs across the entire PHAT footprint. The average stellar density and 50% completeness limits of these ASTs is tabulated in Table 1. However, because of the wide areal coverage of the halo data (), running ASTs for each position of the UKIRT footprint is prohibitively computationally expensive in the framework of this photometric routine. Instead, we run ASTs in three representative stellar density ranges, using stellar density as an indicator of the expected crowding (and resulting photometric noise) in each range. We then measure the average stellar density of the six halo regions discussed in §5.3. We use the set of ASTs that were generated for regions of similar stellar density. Table 2 lists the properties of the ASTs used. The process of generating ASTs in select regions of specific stellar density and applying them more broadly is also used by the PHAT/PHATTER teams, who faced a similar computational challenge of running ASTs across the entire extent of their M31 and M33 footprints, respectively (Williams et al., 2017; Lazzarini et al., 2022). We summarize our ASTs in Table 2.
@@ 5.1 Application to Mock Data
To measure the AGB star SFH, we run MATCH on the ‘observed’ CMDs shown in Figure 4. We adopt an age range of to with a spacing of 0.1 dex and a metallicity range of [M/H] = -1.4 to 0.0 dex with a spacing of 0.2 dex. We use well-populated basis functions () to minimize any effects of Poisson variations, which are the not purpose of this test.
Next, we run the simulated CMDs through MATCH as if they were real photometric datasets. We then compute random uncertainties using HMC. We ran an exhaustive set of simulations with various SFHs. For brevity, we only explore a couple of illustrative examples below for a constant and bursty SFH. However, the conclusions are the same no matter in the input SFH: MATCH self-consistently recovers the input SFH using mock data, within the expected errors.
Synthetic Galaxy I: Multi-burst SFH. Synthetic galaxy I was created from three star formation bursts with ages of 2.0, 5.0, 7.6 Gyr and metallicities [M/H] = -0.4, -0.2, 0.0 dex, respectively. Each burst forms a stellar mass of . In the top panel of Figure 4, we show the “observed” CMD, recovered best fit model CMD, and residual CMD, along with the recovered and input star formation histories.
@@
The excellent agreement between all input and recovered tests indicates that MATCH is self-consistently able to handle our custom basis functions. We now turn to application of this method to real data. We show two further simulated tests in Appendix B for ages up to 14 Gyr and with JWST filters.
@@ 5.2 Comparison to PHAT
A further check on consistency is to compare our AGB star SFHs to those derived from deep optical CMDs. Ideally, we would use an optical CMD that reaches the oldest MSTO to provide an anchor for this testing.
However, the only datasets that are suitably deep, outside the Magellanic Clouds that have been used to calibrate the AGB models, are from small-area HST or JWST fields (e.g., Skillman et al. 2003; Dolphin et al. 2003; Brown et al. 2006; Geha et al. 2015; Albers et al. 2019; McQuinn et al. 2024a) or from ground-based imaging of Milky Way (MW) satellites (e.g., de Boer et al. 2011, 2012a, 2012a, 2014; Muñoz et al. 2018). At issue is that these fields do not contain a suitable number of AGB stars (at least , as we discuss below) to enable a reliable SFH.
We instead opt to compare our AGB star SFH to those measured as part of the PHAT survey. Specifically, we use the SFH fits from W17, who measured the spatially resolved ancient SFH of the PHAT region from optical CMDs that extend to mag below the RC, as shown in Figure 5. Although these CMDs do not reach the 10 Gyr MSTO, and instead rely on advanced phases of evolution (e.g., the position and shape of the RC), W17 took great care to validate their results against CMDs taken throughout M31’s disk that did reach the oldest MSTO (Bernard et al., 2015a, b). We did consider comparing results directly to other M31 SFHs that do reach the oldest MSTO (Brown et al., 2006; Richardson et al., 2008; Bernard et al., 2015a), but there are not enough AGB stars in the spatially matched HST footprints.
W17 measured the SFH of M31’s outer disk up to 14 Gyr. Recognizing the limitations of their moderately deep CMDs, W17 adopt a single ancient age bin of 8-14 Gyr ago. They adopt a finer time resolution (0.1 dex) for ages younger than Gyr. For this study, to probe the oldest AGB-star SFH for ages of Gyr, complete, well-measured photometry of the faintest (and oldest) AGB stars are needed. Unfortunately, because our data were taken for different primary science goals, the photometry of these oldest AGB stars have high uncertainties; the 50% completeness limit of our data that overlaps with the PHAT footprint is at mag. This is approximately where the oldest AGB stars lie in magnitude; all the AGB stars with ages Gyr are fainter than J = 18.3 mag (see Figure 5). Therefore, we adopt J = 18.0 mag as the lower magnitude limit because the largest difference between the injected and recovered ASTs is 0.3 mag at that magnitude, and thus no AGB stars with ages Gyr will be present above J = 18 mag. In conclusion, due to the current limits of the data in this study, we only measure the AGB-based SFHs back to 8 Gyr ago. We stress this age limit is not inherent to the AGB star SFH method itself, but rather due to our desire for empirical validation.
We also limited the fitting routine to follow the chemical evolution model in W17, which involved limiting our fits to only include the metallicities allowed in Figure 8 of W17 for M31’s outer disk (blue line, kpc). In a future paper, we plan to incorporate more sophisticated chemical enrichment modeling schemes in our code.
@@ 5.2.1 Photometric Selection
Beyond an age restriction, we limit this comparison to the outer region ( kpc) of the PHAT survey, as illustrated in Figure 6. This enables us to compare to the deepest optical data available, and avoids the need to model the complex spatially varying differential extinction that dominates most of M31’s inner disk (e.g., Dalcanton et al., 2015; Lewis et al., 2015). It also mitigates the impact of crowding on the optical CMDs and hence the SFHs. Crowding significantly impacts the depth of the optical CMDs in M31’s inner disk, reducing the fidelity of the W17 SFHs. We note that while crowding affects the faintest stars in the the PHAT survey, it has minimal impact on brighter stars. For example, Figure 6 in Gregersen et al. (2015) demonstrates the RGB stars with kpc have completeness in the PHAT survey.
@@
In comparison, our AGB star SFH contains AGB stars that are located within the selection box in Figure 5. Unlike the the PHAT optical CMDs, our AGB star data are not generally crowded in any part of M31. The 50% completeness limit in the inner regions ( kpc) is only mag brighter than in the outskirts. The significantly reduced sensitivity to crowding is another strength of AGB stars in the NIR for SFH determinations.
. We excluded data inside a deprojected distance of kpc, denoted by the shaded grey hatched ellipse, to avoid the crowded and reddened inner regions of M31’s disk that significantly affect the optical CMD.
@@
We opt to use the Draine et al. (2014) dust map, which is based on far-infrared and submillimeter emission observations from the Spitzer Space Telescope and Herschel Space Observatory. The Dalcanton et al. (2015) reddening map predicts less extinction than the Draine et al. (2014) dust map, although their overall morphological agreement is excellent. Thus, we use the Draine et al. (2014) map with the Dalcanton et al. (2015) corrections applied. Specifically, the Dalcanton et al. (2015) corrects the Draine et al. (2014) maps by dividing by a factor of and adding a bias of to account for possible systematic biases. This transformation is then as follows: .
@@
Figure 5 shows a comparison between the HST-based optical CMD of M31 at kpc (left) and our ground-based NIR CMD of the same region (right). As discussed in §2, it is virtually impossible to separate out AGB and RGB stars in the optical. This point is reinforced by the AGB isochrones overlaid on the optical PHAT CMD.
In the NIR, we select AGB stars as having colors between mag and magnitudes between mag based on the locations of AGB stars isochrones, and due to the limited completeness of the fainter, older AGB stars, as discussed in Section 5.2. This photometric selection region also contains red helium-burning stars which will provide additional age information for young ( Myr) populations. These criteria also eliminate contamination from foreground stars, which are bluer than our selection region. We also emphasize that most of the age information in the PHAT SFHs is from the position of the red clump at mag, whereas we only used AGB stars brighter than mag in NIR wavelengths.
5.2.2 Measuring the AGB star SFH in the PHAT Region
@@
To measure the AGB star SFH, we run MATCH on stars within the grey box in the UKIRT CMD shown in Figure 5. For our fitting, we adopt an M31 distance modulus of mag ( kpc; Lee 2023). This distance was derived from the J-region asymptotic giant branch (JAGB) method, a standard candle based on AGB stars, using the same UKIRT photometry studied in this paper, and agrees to within 1 with other commonly adopted distances to M31 measured with the TRGB, RR Lyrae, and Cepheids (e.g., McConnachie et al., 2005; Jeffery et al., 2011; de Grijs & Bono, 2014; Bhardwaj et al., 2016; Savino et al., 2022). We use a CMD bin size of mag in color and magnitude, respectively. We then follow the CMD modeling procedure described in §3.
@@
Figure 7 shows the SFH of the PHAT region based on AGB stars. From the residual CMDs (left panel), we see that our best fit model provides a good match to the data. The residual significance plot, i.e., the residual weighed in each Hess diagram in units of standard deviations, shows no strong artifacts but a rather uniform ‘checkerboard’ pattern. This indicates that there are no regions of the CMD that are systematically discrepant. This plot shows modest discrepancies between the data and model at the bluest colors, but they are of moderate significance and are on-par with the quality of residuals in most optical CMD modeling (e.g., Monelli et al. 2010; Hidalgo et al. 2011; Weisz et al. 2011, 2014; Skillman et al. 2017; Albers et al. 2019; Savino et al. 2023) and with JWST in the NIR (e.g., McQuinn et al., 2024a, c).
@@
The middle panel shows the cumulative SFH the AGB star SFH in black, and the W17 best fit PARSEC-based SFH in red. The uncertainty envelope for both SFHs reflects the 68% random uncertainties. We note that although W17 include AGB stars in their CMD modeling they (a) are of limited utility due to overlap with the RGB and (b) are not fit with the more recent versions of the COLIBRI AGB models, as they were not yet available. As discussed extensively in W17, the age information at most epochs comes from the RC and ratios of RC to RGB stars, with some contribution from AGB stars. The two SFHs (AGB star and W17) use the same IMF, but adopt slightly different minimum mass limits for the low-mass IMF. This introduces an order of unity difference in the SFHs and stellar masses formed. We apply this small IMF correction to the W17 for the purposes of equal comparison.
@@
The two cumulative SFHs are in excellent agreement to within . Both formed the same stellar mass. The AGB star SFH has a burst Gyr ago, while the optical SFH shows a burst of similar amplitude from Gyr ago. Both show a modest amount of stellar mass formed Gyr and Gyr ago. Though the difference in timing is not statistically significant, in future studies it will be interesting to understand if such differences could be due to information content of the data (e.g., ‘bleeding’ of star formation into adjacent time bins due to the increased age-metallicity degeneracy in the RC).
@@
We also compared the age-metallicity relationships (AMRs) from our fits and W17 in Appendix C. Both are generally consistent with solar metallicity, except for a noticeable drop to [M/H]=-0.5 at Gyr. Similar decreases in metallicity are measured in other regions of M31’s disk (Bernard et al., 2015a) from CMDs that reach the oldest MSTO.
@@
The right panel of Figure 7 shows the differential SFH, i.e., SFR vs time, of the AGB star and PHAT SFHs. As with the cumulative SFHs, we find good agreement between the amplitude and timing of star formation in both SFHs, though there are a few differences. For example, the timing and amplitude of star formation around Gyr and Gyr ago are slightly different. As with the cumulative SFHs, such differences in the exact timing of star formation from CMD modeling are often observed when using different stellar models or CMD fitting codes (e.g., Weisz et al., 2011; Dolphin, 2012; Cole et al., 2014; Skillman et al., 2017; Savino et al., 2023).
@@
To examine how significant these variations may be, in Figure 8 we compare best fit SFHs measured using different stellar models from W17 with our AGB star SFH. It is clear from the cumulative and differential SFHs that our AGB star SFH is well within the model-to-model variance from optical CMD analysis. Each of the optical-based SFHs shows the same qualitative trend of stellar mass growth, but with slightly different timings and amplitudes of star formation episodes, as well as modest differences in the total stellar mass formed. Within this context, our AGB-based SFHs are as robust as any of the SFHs based on deeper optical CMDs.
@@
Our measurement is based on AGB stars. In Appendix D, we quantify how the quality of the fit and statistical errors scale with the number of AGB stars, by measuring the AGB star SFH from decreasingly smaller random samples of AGB stars. In Figure D1, we show that as few as 1000 AGB stars can deliver a reasonable (i.e., recovering bursts at roughly the correct time, albeit with much larger uncertainties on the amplitudes of these bursts) SFH compared to the W17.